# The neutrino behavior in the Einstein-Hilbert-Bumblebee gravity around global monopole field

### v1

@@ VI Numerical investigation: two flavor case
$b^{2}\left(\frac{x_{D}}{r_{s}}+1-\frac{y_{D}\pi}{2r_{s}}(a+\overline{\mu})\right)-\frac{4My_{D}b}{(1-\overline{\mu})r_{s}}=\sqrt{1-\left(\frac{b}{r_{s}}\right)^{2}}\left(y_{D}b+(a+\overline{\mu})\frac{x_{D}\pi b}{2}+\frac{4M}{(1-\overline{\mu})}x_{D}\right),$ (25)
here we assume that a circular trajectory of the detector around the Sun (with $x_{D}\ =\ r_{D}\cos{\phi}$, $y_{D}=r_{D}\sin{\phi}$), thus we can numerically deal with Eq.(25) to get the impact parameters.
@@
Next we shall calculate the oscillation probability of neutrinos that are lensed by a Schwarzschild-like black hole in the EHB gravity around global monopole field. In order to have a quantitative understanding of oscillation probability, we consider a Sun-Earth system with the Sun as the lens and the Earth as a detector $(r_{D}=\ {10}^{8}km)$. Assuming that the gravitational field of the Sun is represented by the exterior of the Schwarzschild-like metric in the EHB gravity around global monopole field in the weak field limit, the source is behind the Sun $(r_{s}={10}^{5}r_{D})$ and it emits high-energy neutrinos with $E_{0}=\ 10MeV$. Then we numerically solve Eq. (19) to obtain the flavor transition probability, by using these numerical values we plot the oscillation probability of neutrinos with respect to the azimuth angle $\phi$. The oscillation probability for the two flavor toy model of neutrinos is shown in Fig. 2, 3, 4, 5 and 6 as a function of the azimuth angle $\phi$, our main purpose is the dependence of oscillation probability on global monopole $\overline{\mu}$.
@@
Fig. 2 shows the dependence of oscillation probability on the neutrino mass ordering for the nonmaximal $\gamma$, the solid and the dotted-dashed line correspond to normal hierarchy ($\Delta m^{2}>0$) and inverted hierarchy ($\Delta m^{2}<0$) of neutrinos respectively, which indicates that for the nonmaximal $\gamma$, the probability in the case of inverted hierarchy is large than the normal hierarchy and when the nonmaximal $\gamma$ transitions from $\gamma=\frac{\pi}{6}$ to $\gamma=\frac{\pi}{5}$, the probability of the normal hierarchy increases. It should be noted that the maximal mixing angle is set to be $\gamma=\frac{\pi}{4}$, interestingly, the probability in the case of inverted hierarchy is almost the same as that in the normal hierarchy if $\gamma=\frac{\pi}{4}$, and when $\gamma>\frac{\pi}{4}$, the probability in the case of inverted hierarchy is lower than the normal hierarchy, actually, these behaviors can be verified in Eq.(23). On the other hand, inpired by the works retCASA ; retVilenkin , the selection of the Lorentz-violating parameter $a$ and the global monopole $\overline{\mu}$ are set according to the constraints in some key classical tests, such a selection is to better fit the actual situation of the experiment.
@@
Fig. 3 shows the oscillation probability of gravitationally lensed neutrinos for normal hierarchy and inverted hierarchy in Schwarzschild-like spacetime in the EHB gravity around global monopole field, which indicates that when the Lorentz-violating parameter $a$ is absent, the oscillation probability has a significant deviation from the Schwarzschild case for sufficiently large values of the global monopole $\overline{\mu}$, and while when we consider the Lorentz-violating parameter $a$ and global monopole $\overline{\mu}$ simultaneously, this deviation becomes even more obvious. From the oscillation probability given by Eq.(23) and Fig. 3, we can see that the Lorentz-violating parameter $a$ and the global monopole $\overline{\mu}$ indeed affect the oscillation probability, however, how the two parameters affect the oscillation probability? We depict Fig. 4 and Fig. 5 to explain this problem.
@@
Fig. 5 shows the effect of the Lorentz-violating parameter $a$ on the oscillation probability in the presence of the global monopole $\overline{\mu}$, which indicates that the increasing Lorentz-violating parameter $a$ increases the oscillation probability, and if the global monopole $\overline{\mu}$ is absent, the effect of the Lorentz-violating parameter $a$ on the oscillation probability has been verified in retShi:2025plr .
On the other hand, according to the work retHSW , the frequency of the oscillations of probabilities increases as the lightest neutrino mass, which implies that the lightest neutrino mass also plays a role in regulating the frequency of the oscillations of probabilities. In this line of thought, we compare the oscillation probability of the lightest neutrino with mass and without mass under the background of global monopole field in Fig. 6, it shows that when $\overline{\mu}=10^{-5}$ or $\overline{\mu}=8\times 10^{-5}$, as the lightest neutrino mass increases from 0 to 0.01 eV, the frequency of the oscillations of probabilities will increase, especially when the lightest neutrino mass increases to 0.01 eV, the frequency of the oscillations of probabilities still increases with the increase of the global monopole $\overline{\mu}$ (see the red dashed line and the purple dotted-dashed line), which means that the lightest neutrino mass and the global monopole $\overline{\mu}$ have the same mechanism of influencing the frequency of the oscillations of probabilities, however this also makes it difficult for us to use the neutrino oscillation phenomenon as a messenger to detect global monopole.
@@ VII Conclusion
This work we studied the neutrino behavior in the EHB gravity around global monopole field and discussed how the Lorentz-violating parameter $a$ and the global monopole $\overline{\mu}$ affect neutrino behavior in this curved spacetime. Firstly, we derived the expression of neutrino of probability of oscillation in the EHB gravity around global monopole field by combining the mass shell relation and the expression of neutrino of probability of oscillation in curved spacetime, then in weak field limit, the oscillation probabilities for neutrinos lensed by the gravitational source are obtained, finally, by plotting the probability of $\upsilon_{e}\rightarrow\upsilon_{\mu}$ as a function of the azimuth angle $\phi$ for the normal and inverted ordering of neutrino masses, we qualitatively discussed the effect the geometry of the spacetime of the EBH black holes on the neutrino oscillation probability and compared the oscillation probabilities of neutrinos with the case of lensing by Schwarzschild black hole. Our results show that due to the effect of the Lorentz-violating parameter $a$ and the global monopole $\overline{\mu}$, the curve of the oscillation probability in the EHB gravity around global monopole field and the curve of the oscillation probability in the classic Schwarzschild black hole do not overlap, thus we may view that it offers us a new option by using the gravitational effect of neutrinos to detect celestial bodies, meanwhile, since the neutrino of probability of oscillation in the curved spacetime contains the terms $\sum m_{ij}^{2}$ and $\Delta m_{ij}^{2}$, it is also possible to use the known spacetime geometry to estimate the absolute mass of a single neutrino.
Another important point is that the effects of the Lorentz-violating parameter $a$ and the global monopole $\overline{\mu}$ on probability of oscillation are different. Specifically, as the Lorentz-violating parameter $a$ increases, the peak of the oscillation probability will increase, this obvious difference is expected to meet the current observation accuracy of Icecube-Gen2. Besides, the increase of the global monopole $\overline{\mu}$ and the lightest neutrino mass will both result in an increase for the frequency of the oscillations of probabilities, they have a same mechanism in the terms of the oscillations of probabilities, thus it is a big challege for us to use the neutrino oscillation phenomenon as a messenger to detect global monopole. Fortunately we have other established techniques for observing physical celestial bodies, including the gravitational wave retWu:2022eiv , black hole shadow retWu:2024hxr and so on, they can serve as technical supplements, making the measurement of the global monopole $\overline{\mu}$ possible and improving the detection mechanism for compact celestial bodies.

### v2

@@ VI Numerical results: two flavor case
$b^{2}\left(\frac{x_{D}}{r_{s}}+1-\frac{y_{D}\pi}{2r_{s}}(a+\overline{\mu})\right)-\frac{4My_{D}b}{(1-\overline{\mu})r_{s}}=\sqrt{1-\left(\frac{b}{r_{s}}\right)^{2}}\left(y_{D}b+(a+\overline{\mu})\frac{x_{D}\pi b}{2}+\frac{4M}{(1-\overline{\mu})}x_{D}\right).$ (25)
@@
Next we shall calculate the oscillation probability of neutrinos that are lensed by a Schwarzschild-like black hole in the EHB gravity around global monopole field. In order to have a quantitative understanding of oscillation probability, we consider a Sun-Earth system with the Sun as the lens and the Earth as a detector $(r_{D}=\ {10}^{8}km)$. Assuming that the gravitational field of the Sun is represented by the exterior of the Schwarzschild-like metric in the EHB gravity around global monopole field in the weak field limit, the source is at a distance $(r_{s}={10}^{5}r_{D})$ from the sun and it emits high-energy neutrinos with $E_{0}=\ 10MeV$. Then we assume that the detector follows a circular orbit around the sun ($x_{D}\ =\ r_{D}\cos{\phi}$, $y_{D}=r_{D}\sin{\phi}$), thus by dealing with Eq.(25) numerically to obtain the impact parameters, we can derive the oscillation probability of neutrinos with respect to the azimuth angle $\phi$. The oscillation probability for the two flavor toy model of neutrinos is shown in Fig. 2, 3, 4, 5 and 6 as a function of the azimuth angle $\phi$, our main purpose is the dependence of oscillation probability on global monopole $\overline{\mu}$.
@@
Fig. 2 shows the dependence of oscillation probability on the neutrino mass ordering for the nonmaximal $\gamma$, which indicates that for the nonmaximal $\gamma$, the probability in the case of inverted hierarchy is large than the normal hierarchy and when the nonmaximal $\gamma$ transitions from $\gamma=\frac{\pi}{6}$ to $\gamma=\frac{\pi}{5}$, the probability of the normal hierarchy increases. It should be noted that the maximal mixing angle is set to be $\gamma=\frac{\pi}{4}$, interestingly, the probability in the case of inverted hierarchy is almost the same as that in the normal hierarchy if $\gamma=\frac{\pi}{4}$, and when $\gamma>\frac{\pi}{4}$, the probability in the case of inverted hierarchy is lower than the normal hierarchy, actually, these behaviors can be verified in Eq.(23). On the other hand, inpired by the works retCASA ; retVilenkin ; retPantig:2024kqy , the selection of the Lorentz-violating parameter $a$ and the global monopole $\overline{\mu}$ are set according to the constraints in some key classical tests, such a selection is to better fit the actual situation of the experiment.
@@
Fig. 3 shows the deviation of the oscillation probability of neutrinos in Schwarzschild-like spacetime in the EHB gravity around global monopole field from the case in Schwarzschild black hole, which indicates that when the Lorentz-violating parameter $a$ is absent, the oscillation probability has a significant deviation from the Schwarzschild case for sufficiently large values of the global monopole $\overline{\mu}$, and while when we consider the Lorentz-violating parameter $a$ and global monopole $\overline{\mu}$ simultaneously, this deviation becomes even more obvious. From the oscillation probability given by Eq.(23) and Fig. 3, we can see that the Lorentz-violating parameter $a$ and the global monopole $\overline{\mu}$ indeed affect the oscillation probability, however, how the two parameters affect the oscillation probability? We depict Fig. 4 and Fig. 5 to explain this problem.
@@
Fig. 5 shows the effect of the Lorentz-violating parameter $a$ on the oscillation probability in the presence of the global monopole $\overline{\mu}$, which indicates that the increasing Lorentz-violating parameter $a$ increases the oscillation probability, and if the global monopole $\overline{\mu}$ is absent, the same result has been verified in retShi:2025plr .
On the other hand, according to the work retHSW , the frequency of the oscillations of probabilities increases as the lightest neutrino mass, which implies that the lightest neutrino mass also plays a role in regulating the frequency of the oscillations of probabilities. In this line of thought, we compare the oscillation probability of the lightest neutrino with mass and without mass in the EHB gravity around global monopole field in Fig. 6, it shows that when $\overline{\mu}=10^{-5}$ or $\overline{\mu}=8\times 10^{-5}$, as the lightest neutrino mass increases from 0 to 0.01 eV, the frequency of the oscillations of probabilities will increase, especially when the lightest neutrino mass increases to 0.01 eV, the frequency of the oscillations of probabilities still increases with the increase of the global monopole $\overline{\mu}$ (see the red dashed line and the purple dotted-dashed line), which means that the lightest neutrino mass and the global monopole $\overline{\mu}$ have the same mechanism influencing the frequency of the oscillations of probabilities, however this also makes it difficult for us to use the neutrino oscillation phenomenon as a messenger to detect global monopole.
@@ VII Conclusion
This work we studied the neutrino behavior in the EHB gravity around global monopole field and discussed how the Lorentz-violating parameter $a$ and the global monopole $\overline{\mu}$ affect neutrino behavior in this curved spacetime. Firstly, we derived the expression of the neutrino oscillation phase in the EHB gravity around global monopole field in weak field limit, the oscillation probabilities for neutrinos lensed by the gravitational source are obtained, finally, by plotting the probability of $\upsilon_{e}\rightarrow\upsilon_{\mu}$ as a function of the azimuth angle $\phi$ for the normal and inverted ordering of neutrino masses, we qualitatively discussed the effect of the geometry of the spacetime of the EBH black holes on the neutrino oscillation probability and compared the neutrino oscillation probability with the case of lensing by Schwarzschild black hole. Our results show that due to the effect of the Lorentz-violating parameter $a$ and the global monopole $\overline{\mu}$, the curve of the oscillation probability in the EHB gravity around global monopole field and the curve of the oscillation probability in the classic Schwarzschild black hole do not overlap, thus we may view that it offers us a new option by using the gravitational effect of neutrinos to detect celestial bodies, meanwhile, since the neutrino oscillation probability in the curved spacetime contains the terms $\sum m_{ij}^{2}$ and $\Delta m_{ij}^{2}$, it is also possible to use the known spacetime geometry estimating the absolute mass of a single neutrino.
Another important point is that the effects of the Lorentz-violating parameter $a$ and the global monopole $\overline{\mu}$ on oscillation probability are different. Specifically, as the Lorentz-violating parameter $a$ increases, the peak of the oscillation probability will increase, this obvious difference is expected to meet the current observation accuracy of Icecube-Gen2. Besides, the increase of the global monopole $\overline{\mu}$ and the lightest neutrino mass will both result in an increase for the frequency of the oscillations of probabilities, the effects of them on the oscillation probability are the same, thus it is a big challege for us to use the neutrino oscillation phenomenon as a messenger detecting the global monopole. Fortunately, we have other established techniques for observing physical celestial bodies, including the gravitational wave retWu:2022eiv , black hole shadow retWu:2024hxr and so on, they can serve as technical supplements, making the measurement of the global monopole $\overline{\mu}$ possible and improving the detection mechanism for compact celestial bodies.